OpenStudy (anonymous):

Jobs in the food preparation and serving occupation totaled 2,206,000 in 2000. According to the Bureau of Labor Statistics, jobs in this occupation are scheduled to grow by 30.5% by 2010. How many new jobs will there be in 2010? find the total number of jobs expected in this occupation in 2010

OpenStudy (anonymous):

Ok so same thing... First you have to find what 30.5% is of 2,206,000 this will be represented as x/2,206,000. 30.5%=30.5/100 cross multiply 100(x)=100x 30.5(2,206,000)=67,283,000 now: 67,283,000/100= 672,830 x=672,830

OpenStudy (anonymous):

Now the problem says that jobs are expected to increase so you will add x to 2,206,000: 2,206,000+672,830=2,878,830 Your final answer is: the total number of jobs expected in this occupation in 2010 are 2,878,830 jobs.

OpenStudy (anonymous):

Mmm ok it kinda makes sense. I have like eight more problems like this so ya.

OpenStudy (anonymous):

ha ha well I just answered another one for you:) put another one up and maybe I cna help you understand it better I wish I could draw it out for you to see that would make it so much easier!

OpenStudy (anonymous):

I know its alot more helpful that way since i seem to be more of a visual learner. I havent even looked at my other assignment for math he gave like three pages 80 problems but doing only odd numbers

OpenStudy (anonymous):

Oh ya same here about being a visual learner..... Well i'm willing to help you on some of them

OpenStudy (anonymous):

http://www.dabbleboard.com/draw?b=Guest617698&i=0&c=2a89f602736eb0d2009ee0ad88e0cc975d2b48ee Ok let's try this website real fast go to it and tell me if it is working and if you can see me writing on it....

OpenStudy (anonymous):

the site opened but i dont see you drawing anything at the moment

OpenStudy (anonymous):

hmm ok hold on

OpenStudy (anonymous):

did you get the chat message?